What is the Double declining balance of depreciation (DDB) of four years depreciation schedule for a truck that was purchased for 30000000/= with a life time of 4 years and has residual value of 2000000/=?
1
Expert's answer
2019-10-23T09:15:55-0400

Solution:


We need to find the Double declining balance of depreciation of four years schedule.

We have a formula for Double declining balance of depreciation, that is


"{DDB} = 2 \times Cost \space of \space the \space asset \times Depreciation \space rate"

Here, Cost of the asset = $30000000 and

useful Life = 4 years



"Depreciation \space rate =\frac {1} { Useful \space life} = \frac {1} {4} \space or \space 25 \%"


Depreciation for 1st year

"= 2 \times Cost \space of \space the \space asset \times Depreciation \space rate"



"= 2 \times 30000000 \times \frac {1} {4} = \$15,000,000"

Depreciation for 2nd year

"= 2 \times 15,000,000 \times \frac {1} {4} = \$7,500,000"



Depreciation for 3rd year

"= 2 \times 7,500,000 \times \frac {1} {4} = \$3,750,000"


Depreciation for 4th year

"= 2 \times3,750,000 \times \frac {1} {4} = \$1,875,000"




Answer:


Double declining balance of depreciation (DDB) of four years depreciation = $1,875,000